The Ball Corporation's beverage can manufacturing plant in Fort Atkinson, Wisconsin, uses a metal supplier that provides metal with a known thickness standard deviation σ = .000786 mm. Assume a random sample of 55 sheets of metal resulted in an x¯ = .3307 mm. Calculate the 98 percent confidence interval for the true mean metal thickness. (Round your answers to 4 decimal places.)

The 98% confidence interval is from ____ to _____
Answer #1

We know that the 98% confidence interval for population mean is given by:
Answer #2

Given:

  • Sample mean (xˉ)0.3307mm

  • Population standard deviation (σ)0.000786mm

  • Sample size (n): 55

  • Confidence level: 98%


Step 1: Find the Critical Z-Value (zα/2)

For a 98% confidence interval, the area in each tail is 1%.
Using the standard normal table:

zα/2=z0.012.326

Step 2: Calculate the Margin of Error (E)

E=zα/2×σn=2.326×0.00078655E2.326×0.000106=0.000246

Step 3: Compute the Confidence Interval

Lower bound=xˉE=0.33070.000246=0.3305mmUpper bound=xˉ+E=0.3307+0.000246=0.3309mm

Answer (Rounded to 4 Decimal Places):

The 98% confidence interval for the true mean metal thickness is from

0.3305mmto0.3309mm



answered by: anonymous